Circular No. 8580

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005dd, 2005de, AND 2005df Further to IAUC 8570, E. Lee, N. J. Ponticello, and R. J. Foley report

the LOSS/KAIT discovery of two apparent supernovae on unfiltered KAIT images, as tabulated below. SN 2005de was independently discovered by T. Puckett and J. Tigner at mag 18.0 on an unfiltered CCD frame taken with the 0.60-m patrol telescope (cf. IAUC 8569) on Aug. 3.13 UT.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005dd Aug. 1.46 2^{h}34^{m}36.^{s}57 + 1^{o}20′53′′.2 18.6 10′′.7 E, 1′′.6 N
2005de Aug. 2.28 18 02 23.36 +26 03 07.7 18.1 17′′.3 W, 33′′.1 N

Puckett provides position end figures 23^{s}.37, 07′′.9 for SN 2005de. Additional approximate unfiltered magnitudes from the discoverers: SN 2005dd in UGC 2062, 2004 Jan. 16 UT, [19.0; 2005 Aug. 3.50, 18.8. SN 2005de in UGC 11097, 2002 Aug. 12, [20.0 (Puckett); 2004 Oct. 5, [20.0 (Puckett); 2005 July 28.29, [19.0 (KAIT); Aug. 3.26, 17.6 (KAIT); 4.13, 17.3 (Puckett).

R. Evans, Hazelbrook, N.S.W., reports his visual discovery of an apparent supernova (mag ∼ 13.8) with a 0.31-m reflector on Aug. 4.625 UT. SN 2005df, which is about 15′′ east and 40′′ north of the center of NGC 1559, was not visible on July 29.6 (limiting mag 15.0), and nothing appears at this location to mag ∼ 19 on the European Southern Observatory ‘B’ Survey or on an image of the galaxy in Wassilieff’s CCD Atlas. Evans also found SNe 1984J and 1986L in NGC 1559. A. Gilmore, Mount John Observatory, reports the new object at R = 13.7 on a 10-s CCD image taken with the 0.6-m f/6.4 reflector on Aug. 5.44, providing the following precise position for SN 2005df: α = 4^{h}17^{m}37^{s}.85, δ = −62^{o}46′09′′.5 (equinox 2000.0); nothing is visible at this location on a Digital Sky Survey image.

VARIABLE STAR NEAR IC 1309 Ponticello and Foley also report the discovery of a variable star located

at α = 20^{h}03^{m}00^{s}.24, δ = −17^{o}12′11′′.4 (equinox 2000.0), which is 17′′.5 west and 103′′.4 north of the nucleus of IC 1309. Approximate KAIT magnitudes: July 28.32 UT, [19.0; 29.31, 18.8 (hint of the object); Aug. 2.30, 17.2; 3.31, 17.2. J. Greaves notes that a star is catalogued (in the GSC and USNO-B1.0) near this position, with USNO-A2.0 end figures 00^{s}.271, 12′′.10 (red mag 17.6, blue mag 20.1); he adds that the possible long-period variable is visible on blue and red sky survey plates (1977.5483, 1982.7064, 1988.5462, 1992.3546), but not on 1987.7084 red or 1953.6194 red and blue plates.

2005 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green